2. You are trying to make 1.0 liter a pH 11.5 solution of Ammonia NH3 (Kb= 1.8 *10 -5), how many grams do you need to measure out? (hint, first find the concentration of the solution and then figure out grams.)
1
Expert's answer
2020-06-04T10:37:33-0400

equilibrium reaction involved is;


NH3 + H2O <--->NH4+ + OH-

so at equilibrium we will be having [NH4+]=[OH-]


expressions for Kb = [NH4+][OH-]/[NH3]


so [NH3]=[NH4+][OH-]/Kb

given pH=11.5

so pOH=14-pH= 2.5

[OH-]=10-2.5=[NH4+]

on solving we will get [NH3]=0.55M

so no. of moles of ammonia in 1L= 0.55 moles

mass of ammonia = 0.55*17 = 9.44g